OpenStudy (anonymous):

1/9= 3/5x+ 1/2x+ 3/5/ this is anoter clearing fraction eqation

OpenStudy (anonymous):

\[\frac{1}{9}=\frac{3}{5}x + \frac{1}{2}x + \frac{3}{5}\]\[\frac{1}{9}-\frac{3}{5}=\frac{3}{5}x+\frac{1}{2}x\]\[\frac{5}{45}-\frac{27}{45}=\frac{6}{10}x+\frac{5}{10}x\]\[-\frac{22}{45}=\frac{11x}{10}\]\[x=-\frac{220}{495}=-\frac{4}{9}\]

OpenStudy (anonymous):

man your gud at fractions

jhonyy9 (jhonyy9):

1/9 =3/5x +1/2x +3/5 1/9 -3/5 =3/5x +1/2x 5-27 6 +5 ----- = ------ 45 10x 2x(5-27)=9(6+5) -44x=99 x=99/-44 =-99/44 =-9/4 x=-9/4